Determine mu Subscript x overbar and sigma Subscript x overbar from the given parameters of the population and sample size. muequals72​, sigmaequals24​, nequals36
Answer #1

- NA = Ox=? Ginen w=72,o=24 n=36 MI ECX7372, 02-V(X)=(24)? let X= 14 Exi L ECT)= E(55Xi) = = = E(XC) = KEMD = E(Xi) (Mg = 72